# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ I Introduction
To prepare such complex entangled states using unitary gates that are geometrically local (a natural constraint in many current quantum computing devices), deep circuits are typically required. For instance, any state generated by a logarithmic-depth 1D or constant-depth 2D circuit has sufficiently weak entanglement, such that expectation values of products of local observables can be efficiently computed classically [1]. Despite this simple structure, the problem of sampling from 2D constant-depth circuits—i.e. generating measurement outcomes sampled according to the Born probabilities encoded by the output state—is known to be hard in the worst case [2]. One can gain appreciation for this stark difference by recognising that measurements themselves can convert short-ranged entanglement into long-ranged entanglement—a phenomenon that underpins concepts such as entanglement swapping [3, 4, 5, 6], measurement-based quantum computation [7, 8], and the efficient preparation of topologically ordered states [9, 10, 11, 12, 13, 14]. In the worst-case example mentioned above [2], a macroscopic amount of long-ranged measurement-induced entanglement (MIE) is generated during the sampling process, which is necessary for such physics to be classically hard to simulate.
@@
Based on numerical investigations and non-rigorous analytical arguments, it is frequently conjectured that long-ranged MIE is a typical phenomenon that occurs with high probability upon measuring a subset of sites $B$ in random 2D circuits above some constant critical depth $d_{C}>d_{\rm crit}=O(1)$ [15, 16, 17, 18, 19]. This hypothesised ‘teleportation transition’ [18] from short- to long-ranged MIE as a function of depth is also thought to coincide with a sharp change in the classical hardness of sampling from such circuits [15, 19]. However, due to the nonlinear and stochastic relationship between pre- and post-measurement quantum states, the only rigorous results that prove long-ranged MIE in random circuits so far apply to circuits architectures with qudits whose dimension $q$ grows at least polynomially with system size [20, 19], which, when translated to constant-sized qudits, would require circuits with a depth scaling with system size [21].
@@
In this work, we prove that extensive long-ranged MIE is indeed typical in random 2D circuits of constant depth $d_{C}$ and constant qudit dimension $q$ when a subset of qudits are projectively measured (see Main Result). These include 2D brickwork circuits as well as plaquette circuits made up of 4-local gates (see Fig. 3). By combining our result with previous work relating MIE to the hardness of computing output distributions [15, 19], we provide new evidence that sampling from a random constant-depth 2D circuit yields a quantum advantage: First, the existence of a high-MIE phase implies that standard boundary-contraction strategies based on matrix product states for simulating sampling from a random circuit fail with high probability (Corollary 1). Similarly, boundary-contraction of certain random tensor networks will fail above some critical bond dimension $\chi_{\rm crit}=O(1)$. Second, we prove an unconditional complexity-theoretic separation between random constant-depth 2D circuits and arbitrary sublogarithmic-depth classical circuits, thereby generalizing the worst-case result of Ref. [22] to typical circuit instances.
@@ II Overview of methods and results
This paper focuses on states that can be prepared by two-dimensional geometrically local quantum circuits of constant depth $d_{C}$. Starting from such a state |$\ket{\Psi^{ABC}}$, which we divide into regions $A$, $B$, and $C$, we consider the effect of projective measurements on $B$. Conditioned on a particular measurement outcome $s_{B}$, the unmeasured degrees of freedom collapse into a bipartite post-measurement state |$\ket{\phi^{AC}_{s_{B}}}$ on the unmeasured degrees of freedom $AC$, the entanglement structure of which we wish to study. This setup is illustrated in Fig. 1. We consider several natural families of random circuits, and rigorously prove that extensive long-ranged measurement-induced entanglement is generated in the regime of constant depth $d_{C}=O(1)$. This is shown to have consequences for the hardness of simulating sampling from these circuits, and contracting random tensor networks.
@@
Figure 2: Flow of logic leading to our main technical result, Theorem 8. Blue boxes represent particular quantum states, while yellow boxes with rounded corners represent the key concepts required in our argument. The state |$\ket{\Phi_{d^{\prime}}}$ is a maximally entangled state of dimension $d^{\prime}$, $f(\epsilon)$ is a particular function that vanishes as $\epsilon\rightarrow 0$, and $\mathcal{Z}_{\rm SAW}$ is the partition function of a model of self-avoiding walks. The notation $\widetilde{\log}$ indicates that there are sub-leading corrections to the term $\log\mathcal{Z}_{\rm SAW}$. Most of our intermediate results are very general, applying to any pre-measurement state |$\ket{\Psi}$; the properties of the particular random circuits we consider are only invoked near the end of our proof of Lemma 7 (bottom right). In order to get a succinct expression, a particular assumption about the local measurement bases forming a 2-design is used in Theorem 8 (marked with an asterisk), but we also obtain results for more general setups with minor modifications.
@@
We apply this logic to our setup, with |$\ket{\Psi^{ABC}}$ as the resource state and the measurements on $B$ being in the computational basis; this gives us an entanglement lower bound for the post-measurement states |$\ket{\phi^{AC}_{s_{B}}}$ in terms of the chosen dimension $d^{\prime}$ and protocol error $\epsilon$ (Lemma 6; see the top row of Fig. 2). The problem then becomes quantifying $\epsilon$ as a function of $d^{\prime}$, for this class of entanglement of assistance protocols. In particular, because $B$ is made up of many parties $B=B_{1}B_{2}\ldots$, each of which act locally, and the operations involve only use a single copy of the resource state at a time, this corresponds to a one-shot, multi-user version of entanglement of assistance [34, 35], and our approach leverages tools used to study these kind of problems [27, 28].
@@
Thanks to longstanding rigorous results on these statistical mechanical models [29, 30], the partition function $\mathcal{Z}_{\rm SAW}$ is known to undergo a sharp transition below a critical temperature, where it becomes exponentially suppressed in system size $\mathcal{Z}_{\rm SAW}\leq e^{-\kappa L_{x}}$. The distillation error $\epsilon$ then becomes small even for large values of $d^{\prime}$, and in turn we infer that the MIE is large. In summary, we arrive at the following.
@@
There exist ensembles of unitaries $\mathcal{U}$ obtained from natural circuit architectures on $q$-dimensional qudits of variable depth $d_{C}$, for which we can find values $(q,d_{C})=O(1)$, such that for sufficiently large $L_{x}$, the average measurement-induced entanglement between regions $A$ and $C$ of states generated from random unitaries drawn from this ensemble $U\sim\mathcal{U}$ in the setup shown in Fig. 1 satisfies
@@
for some constant $\kappa>0$.
@@
We now turn to the technical part of the paper, where we formalise the approach described above, state our claims in more detail, and prove the our main theorems.
@@ III.1 Sampling and measurement-induced entanglement
|$\ket{\phi_{s_{B}}^{AC}}=\frac{1}{\sqrt{p_{s_{B}}}}\big{(}\bra{{s_{B}}}\otimes\mathbbm{I}_{AC}\big{)}\ket{\Psi}.$ (3)
@@ III.2 Constant-depth 2D Circuits
Figure 3: Circuit architectures that we explicitly consider in this work. Here, the order of gates runs from bottom to top. (a) A 2D brickwork circuuit, where 2-local gates (blue rectangles) are applied between pairs of qudits in an arrangement that depends on $t\mod 4$, where $t=1,\ldots,d_{C}$ is a discrete time index labeling each layer. (b) 4-local plaquette circuit. By combining groups of 4 two-qudit gates in the 2D brickwork architecture, depth-$2d_{C}$ brickwork circuits can be thought of as special cases of depth-$d_{C}$ 4-local circuits.
@@
While the construction of Ref. [45] demonstrates that all finite-depth states have isoTNS representations, the converse is not true, as the family includes for example the GHZ state and string-net states [45]. We therefore need an extra property which reflects the locality of correlations in these states.
@@ IV.1 Quantifying post-measurement entanglement
$\tilde{\rho}^{A}_{s_{B},U}\coloneqq$ $(\mathbbm{I}_{A}\otimes\bra{s_{B}})\rho^{AB}_{U}(\mathbbm{I}_{A}\otimes\ket{s_{B}})$ (10)
@@ IV.2 Replica-theoretic calculation of quasientropy
$\mathcal{E}^{(2)}_{F:\text{2-design}}=\int{\rm d}\mu_{H}(\phi)\Big{(}\ket{\phi}\!\!\bra{\phi}\Big{)}^{\otimes 2}=\frac{\mathcal{I}_{i}+\mathcal{F}_{i}}{d_{i}(d_{i}+1)},$ (18)
where ${\rm d}\mu_{H}(\phi)$ is the uniform (Haar) measure over all normalized states of dimension $d_{i}$; $\mathcal{I}_{i}=\mathbbm{I}^{\otimes 2}_{Q_{i}}$, and $\mathcal{F}_{i}$ is the swap operator on $Q_{i}\otimes Q_{i}$.
@@
$\mathcal{E}^{(2)}_{B}=\bigotimes_{i\in B}\mathcal{E}^{(2)}_{F_{i}}=\bigotimes_{i\in B}\frac{\mathcal{I}_{i}+\mathcal{F}_{i}}{d_{i}(d_{i}+1)}.$ (20)
@@
$\sum_{s_{B}}q_{s_{B}}\operatorname{Tr}[(\tilde{\rho}^{A}_{s_{B}})^{2}]=$ $\left(\prod_{i\in B}\frac{1}{d_{i}(d_{i}+1)}\right)\mathcal{Z}_{-+}$
@@ V.2 Universal distillation based on random measurements
Since we wish to obtain a bound for arbitrary $d^{\prime}$, we introduce a more general distillation protocol which has an equivalent effect. To specify the channel, we start with a fixed isometry $L_{0}$ from a space $A^{\prime}$ of dimension $d^{\prime}$ to the space $A$ of dimension $d_{A}$. This matrix satisfies $L_{0}^{\dagger}L_{0}=\mathbbm{I}_{A^{\prime}}$, while $L_{0}L_{0}^{\dagger}$ will be a projector on $A$. Now pick an ensemble of unitaries $V_{r_{A}}$ on $A$ with corresponding probabilities $q_{r_{A}}$ that form a unitary 2-design (see e.g. Ref. [63]). The object $L_{r_{A}}^{\dagger}\coloneqq L_{0}^{\dagger}V_{r_{A}}$ then constitutes our random measurement operator. Since the unitaries are drawn from a 2-design, we have that $\sum_{r_{A}}q_{r_{A}}V_{r_{A}}XV_{r_{A}}^{\dagger}=\operatorname{Tr}[X]\pi^{A}$ for all operators $X$, where $\pi^{A}=\mathbbm{I}_{A}/d_{A}$ is the maximally mixed state on $A$. This in turn implies that the map
@@
The next step is to perform local operations on |$\ket{\chi^{A^{\prime}C}_{r_{A}s_{B}}}$ to get as close as possible to |$\ket{\Phi^{A^{\prime}C^{\prime}}_{d^{\prime}}}$. A useful way to quantify how close we can get is to measure how far the reduced states on $A^{\prime}$ are from the maximally mixed state $\pi^{A^{\prime}}=\mathbbm{I}^{A^{\prime}}/d^{\prime}$,
@@
As demonstrated in Ref. [62], based on the value of $\overline{\epsilon}$, one can infer the existence of a local operation on $C$ that converts |$\ket{\chi^{A^{\prime}C}_{r_{A}s_{B}}}$ into state that have an average overlap of at least $1-2\sqrt{\overline{\epsilon}}$ with the maximally entangled state. Thus, $\overline{\epsilon}$ constitutes a figure of merit for the overall EOA task. Because $\overline{\epsilon}$ can be computed without needing to know what the final operation on $C$ actually is, and the channel (26) applied to $A$ does not explicitly depend on $s_{B}$, this quantity will be particularly convenient to work with later on.
@@ V.3 Bounding measurement-induced entanglement in terms of $\overline{\epsilon}$
We take $\rho=\rho^{A^{\prime}}_{r_{A}s_{B}}$ and $\sigma=\pi^{A^{\prime}}$, for which $S(\sigma)=\log d^{\prime}$, which gives $S(\rho^{A^{\prime}}_{r_{A}s_{B}})\geq(1-T)\log d^{\prime}-h_{2}(T)$. Taking the average with respect to $r_{A}$, because $h$ is concave we have $\mathbbm{E}_{r_{A}}h(T)\leq h(\mathbbm{E}_{r_{A}}T)$, and by (28) we have $\mathbbm{E}_{r_{A}}T=\epsilon_{s_{B}}/2$. Together with (33), this yields (31). Applying the same argument to the average over $s_{B}$, $r_{A}$, and $U$ gives Eq. (32). $\square$

### v2

@@ I Introduction
To prepare such complex entangled states using unitary gates alone, deep circuits are typically required. For instance, if we are limited to logarithmic-depth 1D or constant-depth 2D circuits, then the resulting states are simple enough such that expectation values of products of local observables can be efficiently computed classically [1]. It may then seem surprising that the problem of sampling from 2D constant-depth circuits—i.e. generating measurement outcomes sampled according to the Born probabilities encoded by the output state—is known to be hard in the worst case [2]. One can gain appreciation for this stark difference by recognising that measurements themselves can convert short-ranged entanglement into long-ranged entanglement—a phenomenon that underpins concepts such as entanglement swapping [3, 4, 5, 6], measurement-based quantum computation [7, 8], and the efficient preparation of topologically ordered states [9, 10, 11, 12, 13, 14]. In the worst-case example mentioned above [2], a macroscopic amount of long-ranged measurement-induced entanglement (MIE) is generated during the sampling process, which is necessary for such physics to be classically hard to simulate.
@@
Based on numerical investigations and non-rigorous analytical arguments, it is frequently conjectured that long-ranged MIE is a typical phenomenon that occurs with high probability upon measuring a subset of sites in random 2D circuits above some constant critical depth $d_{C}>d_{\rm crit}=O(1)$ [15, 16, 17, 18, 19]. This hypothesised ‘teleportation transition’ [18] from short- to long-ranged MIE as a function of depth is also thought to coincide with a sharp change in the classical hardness of sampling from such circuits [15, 19]. However, due to the nonlinear and stochastic relationship between pre- and post-measurement quantum states, the only rigorous results that prove long-ranged MIE in random circuits so far apply to circuits architectures with qudits whose dimension $q$ grows at least polynomially with system size [20, 19], which, when translated to constant-sized qudits, would require circuits with a depth scaling with system size [21].
@@
In this work, we prove that extensive long-ranged MIE is generated in random 2D circuits of constant depth $d_{C}$ and constant qudit dimension $q$ when a subset of qudits are projectively measured (see Main Result). These include 2D brickwork circuits as well as plaquette circuits made up of 4-local gates (see Fig. 3). By combining our result with previous work relating MIE to the hardness of computing output distributions [15, 19], we provide new evidence that sampling from a random constant-depth 2D circuit yields a quantum advantage: First, the existence of a high-MIE phase implies that standard boundary-contraction strategies based on matrix product states for simulating sampling from a random circuit fail with high probability (Corollary 1). Similarly, boundary-contraction of certain random tensor networks will fail above some critical bond dimension $\chi_{\rm crit}=O(1)$. Second, we prove an unconditional complexity-theoretic separation between random constant-depth 2D circuits and arbitrary sublogarithmic-depth classical circuits (Corollary 2), thereby generalizing the worst-case result of Ref. [22] to typical circuit instances.
@@ II Overview of methods and results
This paper focuses on states that can be prepared by two-dimensional geometrically local quantum circuits $U$ of constant depth $d_{C}$. Starting from such a state |$\ket{\Psi^{ABC}}$, which we divide into regions $A$, $B$, and $C$, we consider the effect of projective measurements on $B$. Conditioned on a particular measurement outcome $s_{B}$, the unmeasured degrees of freedom collapse into a bipartite post-measurement state |$\ket{\phi^{AC}_{s_{B}}}$ on the unmeasured degrees of freedom $AC$, the entanglement structure of which we wish to study. This setup is illustrated in Fig. 1. We consider several natural families of random circuits, and rigorously prove that extensive long-ranged measurement-induced entanglement is generated in the regime of constant depth $d_{C}=O(1)$. This is shown to have consequences for the hardness of simulating sampling from these circuits, and contracting random tensor networks.
@@
Figure 2: Flow of logic leading to our primary technical finding, Theorem 8, and in turn our main result—that the post-measurement states |$\ket{\phi^{AC}_{s_{B}}}$ possess extensive measurement-induced entanglement (scaling with the linear system size $L_{x}$). Top row: After measuring on $B$, we consider the effect of some distillation protocol, whose aim is to generate a state that is close to the maximally entangled state |$\ket{\Phi_{d^{\prime}}}$ of rank $d^{\prime}$. Overall, this procedure constitutes an entanglement of assistance task. Right panel: The self-avoiding walk partition function is known to exhibit an ordering transition driven by the temperature $\beta^{-1}$. Here, $\beta$ is set by the amount of entanglement in the pre-measurement state, which is determined by circuit depth and/or qudit dimension. As a result, we find regimes where we can lower bound MIE in the form (1).
@@
We apply this logic to our setup, with |$\ket{\Psi^{ABC}}$ as the resource state and the measurements on $B$ being in the computational basis (see the top row of Fig. 2); this gives us an entanglement lower bound for the post-measurement states |$\ket{\phi^{AC}_{s_{B}}}$ in terms of the chosen dimension $d^{\prime}$ and protocol error $\epsilon$ (Lemma 6). The problem then becomes quantifying $\epsilon$ as a function of $d^{\prime}$, for this class of entanglement of assistance protocols. In particular, because $B$ is made up of many parties $B=B_{1}B_{2}\ldots$, each of which act locally, and the operations involve only use a single copy of the resource state at a time, this corresponds to a one-shot, multi-user version of entanglement of assistance [34, 35]. As we shall see later (Sec. V.4), it will be possible to leverage certain tools that were developed to study these kind of problems [27, 28].
@@
Thanks to longstanding rigorous results on these statistical mechanical models [29, 30], the partition function $\mathcal{Z}_{\rm SAW}$ is known to undergo a sharp transition below a critical temperature, where it becomes exponentially suppressed in system size $\mathcal{Z}_{\rm SAW}\leq e^{-\kappa L_{x}}$ (see right panel of Fig. 2). The distillation error $\epsilon$ then becomes small even for large values of $d^{\prime}$, and in turn we infer that the MIE is large. In summary, we arrive at the following.
@@
There exist ensembles of unitaries $\mathcal{U}$ obtained from natural 2D circuit architectures on $q$-dimensional qudits of variable depth $d_{C}$, for which there are values $(q,d_{C})=O(1)$, such that the average measurement-induced entanglement between regions $A$ and $C$ in the setup shown in Fig. 1 satisfies
@@
for some constant $\kappa>0$ for sufficiently large $L_{x}$, where $L_{x}$ is the linear system size.
@@
We now turn to the technical part of the paper, where we formalise the approach described above, state our claims in more detail, and prove the main theorems.
@@ III.1 Sampling and measurement-induced entanglement
|$\ket{\phi_{s_{B}}^{AC}}=\frac{1}{\sqrt{p_{s_{B}}}}\big{(}\bra{{s_{B}}}\otimes I_{AC}\big{)}\ket{\Psi}.$ (3)
@@ III.2 Constant-depth 2D Circuits
Figure 3: Circuit architectures that we explicitly consider in this work. Here, the order of gates runs from bottom to top. (a) A 2D brickwork circuit, where 2-local gates (blue rectangles) are applied between pairs of qudits in an arrangement that depends on $t\mod 4$, where $t=1,\ldots,d_{C}$ is a discrete time index labeling each layer. (b) 4-local plaquette circuit. By combining groups of 4 two-qudit gates in the 2D brickwork architecture, depth-$2d_{C}$ brickwork circuits can be thought of as special cases of depth-$d_{C}$ 4-local circuits.
@@
While the construction of Ref. [45] demonstrates that all finite-depth states have isoTNS representations, the converse is not true, as the family includes, for example, the Greenberger-Horne-Zeilinger (GHZ) state and string-net states [45]. We therefore need an extra property which reflects the locality of correlations in these states.
@@ IV.1 Quantifying post-measurement entanglement
$\tilde{\rho}^{A}_{s_{B},U}\coloneqq$ $(I_{A}\otimes\bra{s_{B}})\rho^{AB}_{U}(I_{A}\otimes\ket{s_{B}})$ (10)
@@ IV.2 Replica-theoretic calculation of quasientropy
$\mathcal{E}^{(2)}_{F:\text{2-design}}=d_{i}\int{\rm d}\mu_{H}(\phi)\Big{(}\ket{\phi}\!\!\bra{\phi}\Big{)}^{\otimes 2}=\frac{\mathcal{I}_{i}+\mathcal{F}_{i}}{d_{i}+1},$ (18)
where ${\rm d}\mu_{H}(\phi)$ is the uniform (Haar) measure over all normalized states of dimension $d_{i}$; $\mathcal{I}_{i}=I^{\otimes 2}_{Q_{i}}$, and $\mathcal{F}_{i}$ is the swap operator on $Q_{i}\otimes Q_{i}$.
@@
$\mathcal{E}^{(2)}_{B}=\bigotimes_{i\in B}\mathcal{E}^{(2)}_{F_{i}}=\bigotimes_{i\in B}\frac{\mathcal{I}_{i}+\mathcal{F}_{i}}{d_{i}+1}.$ (20)
@@
$\sum_{s_{B}}q_{s_{B}}\operatorname{Tr}[(\tilde{\rho}^{A}_{s_{B}})^{2}]=$ $\left(\prod_{i\in B}\frac{1}{d_{i}+1}\right)\mathcal{Z}_{-+}$
@@ V.2 Universal distillation based on random measurements
Since we wish to obtain a bound for arbitrary $d^{\prime}$, we introduce a more general distillation protocol which has an equivalent effect. To specify the channel, we start with a fixed isometry $L_{0}$ from a space $A^{\prime}$ of dimension $d^{\prime}$ to the space $A$ of dimension $d_{A}$. This matrix satisfies $L_{0}^{\dagger}L_{0}=I_{A^{\prime}}$, while $L_{0}L_{0}^{\dagger}$ will be a projector on $A$. Now pick an ensemble of unitaries $V_{r_{A}}$ on $A$ with corresponding probabilities $q_{r_{A}}$ that form a unitary 2-design (see e.g. Ref. [63]). The object $L_{r_{A}}^{\dagger}\coloneqq L_{0}^{\dagger}V_{r_{A}}$ then constitutes our random measurement operator. Since the unitaries are drawn from a 2-design, we have that $\sum_{r_{A}}q_{r_{A}}V_{r_{A}}XV_{r_{A}}^{\dagger}=\operatorname{Tr}[X]\pi^{A}$ for all operators $X$, where $\pi^{A}=I_{A}/d_{A}$ is the maximally mixed state on $A$. This in turn implies that the map
@@
The next step is to perform local operations on |$\ket{\chi^{A^{\prime}C}_{r_{A}s_{B}}}$ to get as close as possible to |$\ket{\Phi^{A^{\prime}C^{\prime}}_{d^{\prime}}}$. A useful way to quantify how close we can get is to measure how far the reduced states on $A^{\prime}$ are from the maximally mixed state $\pi^{A^{\prime}}=I^{A^{\prime}}/d^{\prime}$,
@@
As demonstrated in Ref. [62], based on the value of $\overline{\epsilon}$, one can infer the existence of a local operation on $C$ that converts |$\ket{\chi^{A^{\prime}C}_{r_{A}s_{B}}}$ into states that have an average overlap of at least $1-2\sqrt{\overline{\epsilon}}$ with the maximally entangled state. Thus, $\overline{\epsilon}$ constitutes a figure of merit for the overall EOA task. Because $\overline{\epsilon}$ can be computed without needing to know what the final operation on $C$ actually is, and the channel (26) applied to $A$ does not explicitly depend on $s_{B}$, this quantity will be particularly convenient to work with later on.
@@ V.3 Bounding measurement-induced entanglement in terms of $\overline{\epsilon}$
We take $\rho=\rho^{A^{\prime}}_{r_{A}s_{B}}$ and $\sigma=\pi^{A^{\prime}}$, for which $S(\sigma)=\log d^{\prime}$, which gives $S(\rho^{A^{\prime}}_{r_{A}s_{B}})\geq(1-T)\log d^{\prime}-h_{2}(T)$. Taking the average with respect to $r_{A}$, because $h_{2}$ is concave we have $\mathbbm{E}_{r_{A}}h_{2}(T)\leq h_{2}(\mathbbm{E}_{r_{A}}T)$, and by (28) we have $\mathbbm{E}_{r_{A}}T=\epsilon_{s_{B}}/2$. Together with (33), this yields (31). Applying the same argument to the average over $s_{B}$, $r_{A}$, and $U$ gives Eq. (32). $\square$